OpenStudy (anonymous):

2c+3 over C^2-25 plus 7 Over c^2-25

OpenStudy (anonymous):



OpenStudy (campbell_st):

ok so you have a common denominator so add the numerators \[\frac{2c + 3}{c^2 - 25} + \frac{7}{c^2 _ 25} = \frac{ 2c + 3 + 7}{c^2 -25}\]

OpenStudy (campbell_st):

the next part of the solution if to factorise both the numberator and denominator and then eliminate any common factors